OpenStudy (jenniferjuice):

PLEASE HELP???? Tristan says that solutions to radical equations are extraneous. Hadley says that solutions to radical equations are non-extraneous. Find an equation where Tristan is correct and a second equation where Hadley is correct. Using complete sentences, explain each step when solving to justify your examples.

OpenStudy (jenniferjuice):

@IMStuck @iPwnBunnies @iambatman

OpenStudy (jenniferjuice):

@precal

OpenStudy (vishweshshrimali5):

Okay. First of all do you understand the meaning of extraneous solution ?

OpenStudy (jenniferjuice):

yes i do

OpenStudy (vishweshshrimali5):

Great !

OpenStudy (jenniferjuice):

when you solve for x and then when you go to plug it in it does not equal the same solution as it said beofre

OpenStudy (vishweshshrimali5):

So I will give you one example for both persons and you find out another example

OpenStudy (vishweshshrimali5):

\[\large{\sqrt{x} = x-6}\] \[\large{\implies x = (x-6)^2}\] \[\large{\implies x^2 - 13x + 36 = 0}\] \[\large{\implies x^2 - 9x - 4x + 36 = 0}\] \[\large{\implies (x-9)(x-4) = 0}\] \[\large{\implies x = 9,4}\] Though, when you will check you will notice that x = 4 doesn't satisfy the main equation

OpenStudy (vishweshshrimali5):

Extraneous solution ^^

OpenStudy (vishweshshrimali5):

As for non extraneous solution: \[\large{\sqrt{x+3} = 0}\] \[\large{\implies x+3 = 0}\] \[\large{\implies x = -3}\] There is only one solution and this satisfies the main equation ....

OpenStudy (vishweshshrimali5):

@jenniferjuice, did you get these examples ?

OpenStudy (jenniferjuice):

yesssss i diddd

OpenStudy (vishweshshrimali5):

Great can you think of some more ?